OpenStudy (anonymous):

Find the area of a circle that has a radius of 14 feet. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

3.1415926535 * 14^2 = 615.752

OpenStudy (anonymous):

which is roughly 615.75

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the area of a circle is A= π r^2 and if the radius is 14 then your square 14 then multiply it to pi so 14^2 = 196 196 * π = 615.75